News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 2 September 2005 9.00am EDT

During the last week the science program was resumed at 10:15am on 27 Aug following the high radiation event on Aug 22 and the unexpected radiation safing event on Aug 24 that resulted in being LETG left in a partially inserted state. Real-time procedures were completed on Aug 25 to retract the LETG and reset the radiation monitoring (radmon) process to prepare for the resumption of the science mission.

The cause of the unexpected event was determined to be due to a low probability timing condition in the flight software that resulted in the radmon process reading stale data. An operational workaround has been implemented to mitigate future occurrences.

The scheduled targets impacted by the events were PKS 1030-357, DLS J1359-1055, RGB1324+576, Kelu-1, SDSS J233837.10-00, RX J0720.4-3125, IC 1613, X1908+375 and SMCwing 20 (total 249 Ks).

A Chandra image release was issued on 31 August of the star cluster Trumpler 14 showing approximately 1,600 stars and a diffuse glow from hot multimillion degree X-ray producing gas. For details see http://chandra.harvard.edu/photo/2005/trump/

The schedule of targets for last week and next week is shown below.

------------------------------------------
2PIGGz0.061 J0011       ACIS-I       Aug 28
SDSS J005009.81-00      ACIS-S
EGS-7                   ACIS-I
tnj1338-194             ACIS-S
Radiation Belts                      Aug 29
AE Aqr                  ACIS-S/HETG  Aug 30
EGS-7                   ACIS-I       Aug 31
tnj1338-194             ACIS-S
Radiation Belts
ArLac (21 obs)          HRC-S        Sep 1
cl0159+0030             ACIS-I       Sep 2
cl1202+5751             ACIS-I
EGS-7                   ACIS-I       Sep 3
tnj1338-194             ACIS-S
Radiation Belts
Puppis A NE Shock       ACIS-I       Sep 4
IC 1613                 ACIS-S
HD93250                 ACIS-S/HETG  Sep 5 
TT Ari                  ACIS-S/HETG
Radiation Belts                      Sep 6
Cygnus A                ACIS-I       Sep 7
cl0216-1747             ACIS-I
PKS 0920-397            ACIS-S       Sep 8
RGBJ0112+3818           ACIS-S
Radiation Belts                      Sep 9
Limiting Window         ACIS-I
The Crab Nebula         ACIS-S/HETG
TT Ari                  ACIS-S/HETG  Sep 10
PKS2155-304             ACIS-S/LETG
RX J0720.4-3125         HRC-S/LETG   Sep 11

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports